OpenStudy (anonymous):

What is the solution of 2(2x + 3) = -3(x + 5)?

OpenStudy (uri):

2(2x + 3) = -3(x + 5) 4x +6 =-3x - 15 4x+3x =-15 -6 7x =-21 x =-21/7 divide and get the answer :P

OpenStudy (anonymous):

Thank You (;

OpenStudy (uri):

yw :)